### v1

@@ 21      Display the original image and attention maps
Spatial Relevance: The spatial relevance can give insights into which parts of the image the model finds important for differentiating between classes. The observations from the Fig 5 are as follow:
  • [label=–]
  • Layer 1: The model’s attention is broadly spread with higher intensity in some central regions. This indicates the model initially focuses on larger areas, possibly capturing the general structure of the image.
  • Layer 2-3: Attention starts to become more localized, with specific regions of high attention emerging. The model begins to identify and focus on more distinct features.
  • Layer 4-7: Attention further refines, highlighting specific areas more sharply. These layers likely focus on intermediate features that are critical for further processing.
  • Layer 8-12: Attention becomes very concentrated on specific, smaller regions. The model is likely focusing on fine-grained details and specific objects or patterns.
In the initial layers, the model’s attention is broadly spread, capturing the overall structure of the image. As we move to higher layers, the attention becomes more localized, focusing on specific features and objects.
@@ 3.2 Performance of ViT+SVM
The pre-trained ViT model weights were sourced from a specified path and loaded into an initialized ViT model. We used the ViT-B/16 model, whose architecture was previously discussed, and obtained it from torchvision.models. The model’s weights were updated using torch.load and subsequently, the model was set to evaluation mode to facilitate feature extraction. Data transformations were defined to standardize the input images. Each image was resized to 224x224 pixels and converted into a tensor. Data loaders were created for both training and testing datasets , with a batch size of 32. The feature extraction process was encapsulated in a specific function. This function is called twice, once for the training data loader and once for the test data loader. This extracts features from each image in the respective datasets and saves them along with their corresponding labels to separate CSV (comma-separated values) files. Feature extraction was performed in a no-gradient context to optimize performance. The parameters used are mention in Table 1.
@@
We extracted features from the ViT in two scenarios: firstly, using the pre-trained weights of ViT-B/16, which were trained on ImageNet; and secondly, by training the ViT from scratch as detailed in Section 3, starting with the ImageNet initial weights up to epoch 8. These extracted features were separately fed as inputs to an SVM to evaluate performance in both scenarios. We conducted 5-fold cross-validation to determine the ROC curve, mean AUC, and mean accuracy. Additionally, we examined both cases using the specified dataset provided in the original data repository.
Table 2: Accuracy and AUC
@@
ViT+SVM (Pre-trained Weights) 0.89 0.88 0.95
ViT+SVM (Transfer Learning: Epoch 8) 0.89 0.95 0.96
Table LABEL:table:k_fold shows the accuracy and AUC (Area Under the Curve) of approaches: ViT+SVM (Pre-trained Weights) and ViT+SVM (Optimal Weights). The models were evaluated on a specified test set and using 5-fold cross-validation. ViT+SVM (Optimal Weights) outperforms ViT+SVM (Pre-trained Weights) in terms of both mean accuracy (0.95 vs. 0.88) and mean AUC (0.96 vs. 0.95) with 5-fold cross-validation. Though, both the approaches achieved the same accuracy on the specified test set, but ViT+SVM (Optimal Weights) generalizes better to unseen data and suggests potentially less over-fitting as evidenced by its higher mean accuracy and AUC with 5-fold cross-validation. Upon examining the mean RoC curves given in Fig 6, it is evident that ViT+SVM (Optimal Weights) is performing consistently as the folds have a small variance in AUC (0.01) and each fold has a high AUC ($>0.95$). On the other hand, ViT+SVM (Pre-trained Weights) have a mean AUC of 0.95 and mean variance of 0.02.
(a) ViT+SVM with Pre-Trained Weights
(b) ViT+SVM with Optimal Weights
@@ 3.3 Comparison of the Models
Table 3 presents the performance metrics of various models on a specific dataset. The comparison shows that each model has its strengths and is suitable for different scenarios.
Performance on Specified test set: ViT with Transfer Learning consistently performs best across most metrics, making it the most balanced and robust model for this task. ViT+SVM (Optimal Weights) excels in precision for the stressed class and recall for the healthy class, making it ideal where high precision in detecting stressed instances and high recall in identifying healthy instances are crucial. ViT (Pre-Trained Weights) provides a balanced performance but does not outperform the other models in any specific metric.
Table 3: Performance of the models for the specified test set.
@@
ViT+SVM (Pre-Trained Weights) 0.92 0.91 0.91 0.84 0.86 0.85 0.8899
ViT+SVM (Optimal Weights) 0.98 0.86 0.91 0.78 0.97 0.87 0.8943
ViT with Transfer Learning 0.95 0.94 0.95 0.89 0.91 0.90 0.9295
K-fold cross validation: The ViT with Transfer Learning model achieved an average accuracy of 86.8%, with F1-scores of 85.6% for the ”Healthy” class and 88.8% for the ”Stressed” class. The model demonstrated consistent performance across the five folds, although there was a slight dip in accuracy in the fifth fold (83%).
The ViT with SVM model achieved a significantly higher average accuracy of 95%, with F1-scores of 93.6% for Class 0.0 and 95.6% for Class 1.0. This model consistently performed well across all folds, with the lowest accuracy being 92.7% and the highest reaching 95.8%.
@@
ViT with Transfer Learning 0.856 0.888 0.868
ViT+SVM with Optimal Weights 0.936 0.956 0.950
The comparative results, as shown in Fig 4 indicate that the ViT with SVM model outperforms the ViT with Transfer Learning model in terms of both accuracy and F1-scores. The average accuracy of the ViT with SVM model is substantially higher (95% compared to 86.8%). Furthermore, the F1-scores for both classes are also higher in the ViT with SVM model, indicating better classification performance for both ”Healthy” and ”Stressed” classes.
The five-fold accuracies for the ViT with SVM model show less variability and higher consistency compared to the ViT with Transfer Learning model, suggesting that the integration of SVM helps in stabilizing the model’s performance across different data splits.
In conclusion, integrating Support Vector Machine with Vision Transformer (ViT) significantly enhances the model’s performance compared to using transfer learning alone. The ViT with SVM model demonstrates superior accuracy and F1-scores, making it a more reliable choice for classification tasks. The higher consistency across folds further reinforces the robustness of the ViT with SVM model. This comparative analysis highlights the benefits of combining advanced deep learning architectures with classical machine learning techniques for improved performance in identifying drought stress.
@@ 4 Conclusion
Our study successfully addressed the challenge of stress identification using smaller datasets by harnessing the feature extraction capabilities of Vision Transformers combined with the classification prowess of SVM. Unlike conventional CNN architectures, Vision Transformers utilize self-attention mechanisms to effectively capture relationships across different parts of the image, enabling them to model long-range dependencies, which is particularly advantageous for complex datasets where traditional CNNs may struggle to capture global context. Moreover, Vision Transformers can handle images of varying resolutions without necessitating architectural modifications, enhancing their flexibility in handling diverse datasets. Our framework also emphasized explainability by generating attention maps, which provide insights into the model’s focus areas within the images, thereby offering transparency in its decision-making process.
However, our comparative analysis revealed that integrating SVM with ViT significantly enhances classification performance. The ViT+SVM model achieved a higher average accuracy of 95%, with F1-scores of 93.6% for stressed images and 95.6% for healthy images, outperforming the ViT with transfer learning approach. This integration not only improved accuracy but also demonstrated greater consistency across different data splits, highlighting the robustness of the ViT+SVM model.
In summary, our study illustrates the superior performance of combining Vision Transformers with Support Vector Machines over traditional transfer learning methods. This approach provides a more reliable and accurate solution for drought stress identification, while the attention maps offer valuable insights into the model’s decision-making process. The findings affirm the potential of advanced deep learning and machine learning techniques in enhancing agricultural practices and decision-making, paving the way for improved crop management strategies.

### v2

@@ 21      Display the original image and attention maps
Spatial Relevance: The spatial relevance can give insights into which parts of the image the model finds important for differentiating between classes. In Layer 1, the model’s attention is broadly distributed with some central intensity, indicating that the model initially captures coarse, global structures of the image. Moving to Layers 2 ,3 and 4, attention becomes increasingly localized, suggesting the model is beginning to identify distinct regions and features relevant to the classification task. Between Layers 5 and 6, the attention patterns grow sharper and more discriminative. These layers appear to focus on intermediate-level features—regions that are potentially indicative of stress patterns but not yet fine-grained. Layers 7 through 11 show a high concentration of attention in specific, small regions with strong contrast, indicating that the model is now attending to fine details, such as localized stress indicators in the vegetation. Notably, Layer 12 diverges from the preceding layers. The attention becomes more diffused and less sharply defined compared to Layers 8–11. This indicates that the final encoder block form a more holistic representation by aggregating information from previous layers, striking a baalnce of local details with global context for the final classification decision.
@@ 3.2 Performance of ViT+SVM
We employed the Vision Transformer (ViT) models for feature extraction and subsequent evaluation through a Support Vector Machine (SVM) classifier. Primarily, we utilized the ViT-B/16 architecture, sourced from torchvision.models, and loaded pre-trained weights using torch.load from a specified path. The model was initialized and set to evaluation mode to facilitate inference-based feature extraction. To standardize the inputs, all images were resized to 224$\times$224 pixels and transformed into tensors. Data loaders were constructed for both the training and testing datasets, using a batch size of 32. Feature extraction was encapsulated in a dedicated function invoked separately for the training and testing data loaders. The resulting features, along with their corresponding labels, were saved to CSV files for downstream processing. This process was executed within a no-gradient context torch.no_grad() to optimize computational efficiency. The hyperparameters and implementation details used for ViT-based feature extraction and classification are summarized in Table 4.
@@
To analyze the influence of model capacity, we extracted features under two experimental scenarios. In the first, we used the ViT-B/16 model (final weights from scenario 8). In the second scenario, we utilized ViT-L/16, taking into account the weights in scenario 4. In both settings, the extracted features were fed into a Support Vector Machine (SVM) for classification. Performance evaluation was conducted on specified test set as well as using 5-fold cross-validation to generate the Receiver Operating Characteristic (ROC) curves, compute the mean Area Under the Curve (AUC), and estimate the mean classification accuracy.
Table 5: Confusion matrix components and test accuracy across for ViT+SVM.
Scenario TP TN FP FN Test Accuracy
ViT-B/16 +SVM 650 353 48 84 0.8837
ViT-L/16 +SVM 640 364 37 94 0.8845
Table 6: Mean Accuracy and AUC for k-fold cross validation
@@
ViT-B/16 +SVM 0.8837 0.9435 0.98
ViT-L/16 +SVM 0.8845 0.9208 0.96
The confusion matrix components, as shown in Table 5 , show that ViT-B/16 produces a lower number of false negatives (FN), a critical metric in drought stress identification tasks, where missing stressed cases can lead to substantial consequences. The overall performance evaluation of the proposed models reveals the superiority of the ViT-B/16 + SVM architecture over its larger counterpart, ViT-L/16 + SVM. Table LABEL:table:k_fold detailed that ViT-B/16 consistently achieved higher mean accuracy, mean AUC, indicating better generalization capability and robustness. Additionally, as shown in Fig. 6(a), the ROC curve for ViT-B/16 exhibits a higher mean AUC (0.98 ± 0.01) compared to ViT-L/16 (0.96 ± 0.01 in Fig. 6(b)), with consistently strong performance across all folds. The tighter clustering of the ROC curves around the upper-left corner for ViT-B/16 also indicates more reliable classification across varying thresholds. These results demonstrate that despite having a smaller architecture, ViT-B/16 offers more accurate and explainable performance, making it a preferable feature extractor for efficient and dependable drought stress detection.
(a) ViT-B/16 +SVM
(b) ViT-L/16 +SVM
@@ 3.3 Comparison of the Models
Performance on Specified test set:
Our two proposed approaches are compared with a previously published model [18] on the same dataset. Table 7 presents a comparative evaluation of the three models for drought stress detection, reporting precision, recall, F1-score, and overall accuracy on the designated test set. The CNN-based framework demonstrates competitive performance, particularly in terms of stressed plant precision (0.9673). However, it lags behind in F1-score and overall accuracy when compared to the Vision Transformer with Transfer Learning. The ViT+SVM model shows relatively weaker performance, especially in classifying healthy plants, suggesting that the SVM integration may not fully exploit the representational capacity of ViTs for this task. The Vision Transformer with Transfer Learning outperforms both the CNN-based and ViT+SVM models across all metrics, making it the most effective model for distinguishing drought-stressed from healthy plants.
Fig. 7 presents the confusion matrices for three models—CNN Framework, ViT+SVM, and ViT with Transfer Learning—which correspond to the performance metrics previously computed and detailed in Table 7 for the specified test set. In drought stress detection, minimizing false negatives (FN) is crucial, as higher FN could lead to significant oversight in applications like drought stress detection, where failing to identify stressed plants might delay critical interventions. Among the three models compared, the Vision Transformer with Transfer Learning (ViT-TL) demonstrates the lowest number of false negatives (73), highlighting its superior sensitivity to actual stress conditions. In contrast, the CNN-based framework and the ViT+SVM model yield higher false negatives—83 and 84 respectively—indicating a comparatively weaker ability to detect all truly stressed plants. Apparently, false positives (FP) are generally less severe than false negatives, they still represent inefficiencies in resource utilization, such as unnecessary irrigation or pesticide application to healthy plants. Both ViT-TL and the CNN framework exhibit equally low false positives (22), whereas the ViT+SVM model produces a significantly higher FP count of 48, suggesting an over-prediction of stress in healthy plants. Overall, ViT-TL achieves a well-balanced performance with the lowest FN and equally low FP, indicating its robustness in both detecting stressed plants accurately and avoiding unwarranted false alarms.
Table 7: Performance comparison of the models for the specified test set.
@@
CNN Based Framework [18] 0.9673 0.8869 0.9252 0.8203 0.9451 0.8788 0.9075
ViT+SVM 0.9312 0.8856 0.9078 0.8078 0.8803 0.8421 0.8837
ViT-TL 0.9678 0.9005 0.9328 0.8385 0.9451 0.8883 0.9162
(a) CNN Framework [18]
(b) ViT+SVM
(c) ViT-TL
Figure 7: Confusion matrices comparison for CNN, ViT+SVM, and ViT with Transfer Learning models.
K-fold cross validation: Table 8 presents a comparative performance analysis between two model configurations: the Vision Transformer (ViT) with transfer learning and ViT-B/16 combined with a Support Vector Machine (SVM) classifier. In the first approach, the ViT model is fine-tuned end-to-end on the target dataset, whereas in the second, ViT is used as a feature extractor, and an SVM classifier is trained on the extracted features. The evaluation was performed using k-fold cross-validation, and the results are reported in terms of the mean F1-scores for both the Healthy and Stressed classes, along with the overall mean accuracy. The ViT with transfer learning demonstrated superior performance with F1-scores of 0.97 and 0.98 for the Healthy and Stressed classes, respectively, and a mean accuracy of 97.43%. In comparison, the ViT-B/16 + SVM configuration achieved F1-scores of 0.93 (Healthy) and 0.95 (Stressed), and a mean accuracy of 94.35%. These results highlight the effectiveness of end-to-end fine-tuning, which enables the model to learn task-specific representations more effectively than the frozen feature extractor approach.
@@
ViT with Transfer Learning 0.97 0.98 0.9743
ViT-B/16 + SVM 0.93 0.95 0.9435
@@ 4 Conclusion
Our study successfully addressed the challenge of stress identification using smaller datasets by harnessing the feature extraction capabilities of Vision Transformers. Unlike conventional CNN architectures, Vision Transformers utilize self-attention mechanisms to effectively capture relationships across different parts of the image, enabling them to model long-range dependencies, which is particularly advantageous for complex datasets where traditional CNNs may struggle to capture global context. Moreover, Vision Transformers can handle images of varying resolutions without necessitating architectural modifications, enhancing their flexibility in handling diverse datasets. Our framework also emphasized explainability by generating attention maps, which provide insights into the model’s focus areas within the images, thereby offering transparency in its decision-making process.
The comparative analysis of a CNN-based framework, ViT-TL, and ViT+SVM showed that Vision Transformers (ViT) with transfer learning significantly improve classification performance. Its performance is consistent across various data splits, thus offering a reliable and accurate solution for drought stress identification, with attention maps providing valuable insights into the model’s decision-making process. These findings highlight the potential of advanced deep learning techniques to enhance agricultural practices and decision-making, paving the way for more effective crop management strategies.
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.
Author Contributions
Aswini Kumar Patra contributed to the study conception and design. He was also responsible for material preparation, data collection, and analysis. The first draft of the manuscript was written by Aswini Kumar Patra, and all authors reviewed and provided feedback on previous versions. Ankit Varshney contributed to the conceptualization of the ViT and SVM combination. Lingaraj Sahoo supervised the research and edited the manuscript. All authors read and approved the final version of the manuscript.
Data Availability
The data used in this study are derived from a publicly available dataset, which can be accessed at https://www.webpages.uidaho.edu/vakanski/Multispectral_Images_Dataset.html. Additionally, the generated data supporting the findings of this study are available from the corresponding author upon reasonable request.